How do you solve this algebra equation?
9k + 6 = 8k + 13

Found 3 solutions by josgarithmetic, ikleyn, MathTherapy:
Answer by josgarithmetic(39618)

Addition Property of Equality and Additive Inverse Property.

.(supply the step)
.(supply the step)
k=7



That can be done in one step or two steps.
That was the response. Use the properties.


----
My earlier "answer" of 19 was wrong. k=7 is the true answer. I did the steps in my head, and did not catch my mistake while handling these steps.

Answer by ikleyn(52788)

.
Distract 8k from both sides.

You will get

k + 6 = 13.

Distract 6 from both sides. You will get

k = 13 - 6 = 7.

Answer. k = 7.


Answer by MathTherapy(10552)


How do you solve this algebra equation?
9k + 6 = 8k + 13
Subtract 8k from both sides
Subtract 6 from both sides
This should give you the value for k, after combination of the like-terms on each side